20) Arm, Inc., has 10,000 shares of 6%, $100 par value, noncumulative preferred stock and 100,000 shares of $1 par value common stock outstanding at December 31, 2013. If the board of directors declares a $200,000 dividend, the
A) preferred stockholders will receive 1/10th of what the common stockholders will receive.
B) preferred stockholders will receive the entire $200,000. (wrong)
C) $60,000 will be held as restricted retained earnings and paid out at some future date.
D) preferred stockholders will receive $60,000 and the common stockholders will receive $140,000.
Answer #1

Dividend for preferred stockholders

= 10,000*100*6% = 60,000

Dividend for common stockholders

= 200,000 - 60,000

= 140,000

Option D is the answer